Each of the following numbers represents a signed decimal number in the 2’s complement system. Determine the decimal value in each case
 10011001
 11101
 01111011
1
Expert's answer
2020-10-12T18:22:32-0400

10011001 = -(1 + 8 + 16) = -25

11101= -(8+4+1) = -13

01111011 = 1+2+8+16+32+64 = 123